Liabilities and Equity Assets 1,700 Certificate of deposit 10,100 Equity Cash $10,700 1,100 Bond $11,800 Total liabilities anPlease help with "D" only... and please slow calculations as I want to learn, not just get answers ;=) .
Answer #1

Solution :- The market value of the equity would be higher because the value of the bond would be higher and the value of the CD would remain unchanged

Now i explain very easily

If the Interest rates decrease in the market than the required rate of return of the persons also decreased as expectations based on market analysis

And we know the value of stock is calculated by sum of the present value of dividends and if required return decrease then automatically Present value Increase

As if Interest rate is 10% then present Value of $100 received after 1 year = 100 / ( 1 + 0.10) = 90.909

Now if interest rate is 20% then present value = 100 / ( 1 + 0.20) = 83.33

Similiarly If interest rate decrease then Present value increase means Price of Equity increase

IF there is any doubt please ask in comments